Ming Chen began a professional practice on June 1 and plans to prepare financial statements at the end of each month. During June, Ming Chen (the owner) completed these transactions: a. Owner invested $60,000 cash in the company along with equipment that had a $15,000 market value b. The company paid $1,500 cash for rent of office space for the month. c. The company purchased $10,000 of additional equipment on credit (payment due within 30 days). d. The company completed work for a client and immediately collected the $2,500 cash earned e. The company completed work for a client and sent a bill for $8,000 to be received within 30 days. f. The company purchased additional equipment for $6,000 cash g. The company paid an assistant $3,000 cash as wages for the month. h. The company collected $5,000 cash as a partial payment for the amount owed by the client in transaction e i. The company paid $10,000 cash to settle the liability created in transaction c. Owner withdrew $1,000 cash from the company for personal use. Required Enter the impact of each transaction on individual items of the accounting equation. (Enter decreases to account balances with a minus sign.)Accounts Equipmen Receivable Accounts Payable M. Chen,_ Capital M. Chen,Revenue Withdrawals Cash Expenses Bal Bal Bal Bal Bal Bal Bal Bal Bal
Answer #1
Cash + Accounts Receivable + Equipment = Accounts
Payable
+ M. Chen, Capital - M. Chen, Withdrawals + Revenue - Expenses
a. 60000 + + 15000 = + 75000 - + -
b. -1500 + + = + - + - 1500
Bal. 58500 + + 15000 = + 75000 - + - 1500
c. + + 10000 = 10000 + - + -
Bal. 58500 + + 25000 = 10000 + 75000 - + - 1500
d. 2500 + + = + - + 2500 -
Bal. 61000 + + 25000 = 10000 + 75000 - + 2500 - 1500
e. + 8000 + = + - + 8000 -
Bal. 61000 + 8000 + 25000 = 10000 + 75000 - + 10500 - 1500
f. -6000 + + 6000 = + - + -
Bal. 55000 + 8000 + 31000 = 10000 + 75000 - + 10500 - 1500
g. -3000 + + = + - + - 3000
Bal. 52000 + 8000 + 31000 = 10000 + 75000 - + 10500 - 4500
h. 5000 + -5000 + = + - + -
Bal. 57000 + 3000 + 31000 = 10000 + 75000 - + 10500 - 4500
i. -10000 + + = -10000 + - + -
Bal. 47000 + 3000 + 31000 = 0 + 75000 - + 10500 - 4500
j. -1000 + + = + - 1000 + -
Bal. 46000 + 3000 + 31000 = 0 + 75000 - 1000 + 10500 - 4500
Answer #2
Ming Chen began a professional practice on June 1 and plans to prepare financial statements At the end of each month. During June, Ming Chen (the owner) completed these transactions: a. Owner invested $60,000 cash in the company along with equipment that had a $15,000 Market value. b. The company paid $1,500 cash for rent of office space for the month. c. The company purchased $10,000 of additional equipment on credit (payment due within 30 Days). d. The company completed work for a client and immediately collected the $2,500 cash earned. e. The company completed work for a client and sent a bill for $8,000 to be received within 30 Days. f. The company purchased additional equipment for $6,000 cash. g. The company paid an assistant $3,000 cash as wages for the month. h. The company Collected $5,000 cash as a partial payment for the amount owed by the client in transaction e. i. The company paid $10,000 cash to settle the liability created in transaction c. j. Owner Withdrew $1,000 cash from the company for personal use. Required (a) Prepare a tabular analysis